# Language Integration in Fine-Tuning Multimodal Large Language Models for Image-Based Regression

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) show promise for image-based regression tasks, but current approaches face key limitations. Recent methods fine-tune MLLMs using preset output vocabularies and generic task-level prompts (e.g., ”How would you rate this image?”), assuming this mimics human rating behavior. Our analysis reveals these approaches provide no benefit over image-only training. Models using preset vocabularies and generic prompts perform equivalently to image-only models, failing to leverage semantic understanding from textual input. We propose Regression via Transformer-Based Classification (RvTC), which replaces vocabulary-constrained classification with a flexible bin-based approach. Unlike approaches that address discretization errors through complex distributional modeling, RvTC eliminates manual vocabulary crafting through straightforward bin increase, achieving state-of-the-art performance on four image assessment datasets using only images. More importantly, we demonstrate that data-specific prompts dramatically improve performance. Unlike generic task descriptions, prompts containing semantic information about specific images enable MLLMs to leverage cross-modal understanding. On the AVA dataset, adding challenge titles to prompts improves correlations from 0.83 to 0.90, a new state-of-the-art. We demonstrate through empirical evidence from the AVA and AGIQA-3k datasets that MLLMs benefit from semantic prompt information surpassing mere statistical biases. This underscores the importance of incorporating meaningful textual context in multimodal regression tasks.
@@ 1 Introduction
Vision-language models trained on massive unlabeled image-language datasets have demonstrated remarkable capacity to extract universal image features, with CLIP [15] achieving impressive zero-shot classification performance on benchmark datasets such as ImageNet. Building on these foundations, Multimodal Large Language Models (MLLMs) have evolved to seamlessly fuse image and text embeddings with generative language capabilities [5, 11, 22]. This has sparked growing interest in transferring MLLM capabilities to image-based regression tasks, including Image Quality Assessment (IQA), Image Aesthetics Assessment (IAA) [4, 20], and AI-Generated Image Quality Assessment (AIGIQA) [21, 14].
@@
Table 1: Comparison of models on the AVA dataset. RvTC achieves state-of-the-art correlation of 0.83 without using any textual input. When data-specific prompts (challenge titles) are added (RvTC+), performance increases to 0.90. “LP” refers to linear probing of the regression head; “+” indicates inclusion of challenge titles in prompts.
@@
2. Enhancing MLLM performance through data-specific prompts: We demonstrate that generic task prompts (e.g., ”How would you rate this image?”) fail to leverage cross-modal capabilities. In contrast, fine-tuning with data-specific prompts yields substantial improvements in correlation results. For example, the prominent IAA dataset AVA includes challenge titles that characterize images with descriptive phrases such as ”Rule of Thirds” or ”Outdoor Macro Shot.” Incorporating these semantic descriptors during fine-tuning improves correlation results from our already state-of-the-art 0.83 to 0.90.
@@ 2 Related work
We highlight the following notable prior works that inform our method.
REgression using CLAssification (RECLA). RECLA involves transforming a regression problem into a classification problem, leveraging the strengths of classification algorithms for predictive tasks [17]. In RECLA, the continuous target variable is discretized into a set of classes, also referred to as bins, effectively categorizing the range of possible values. A classification model is then trained to predict the appropriate bin for each input, and the prediction is often mapped back to a continuous value by a weighted average of the bins’ midpoints. In this work, we show that RECLA seamlessly integrates with pre-trained Multimodal Large Language Models (MLLMs). This integration is robust and a randomly initialized RECLA head can be fine-tuned together with the MLLM without any special modification to the downstream training procedure.
Multimodal Large Language Models. Vision-language models extend the capabilities of traditional language models to incorporate and process multiple modalities, such as image, audio, and video. A key component of Vision-language models is their ability to align representations across modalities, allowing them to understand the relationships between text and other forms of data.
MLLMs, such as LLaVA [11] and mPLUG-Owl2 [22], further enhance this capability by seamlessly fusing image and text embeddings, enabling more complex multimodal reasoning and generation tasks. The architecture of these models involves a combination of modality-specific encoders and a shared transformer-based network that processes the combined representations. The transformer architecture allows the model to capture long-range dependencies and complex interactions between modalities.
@@
Image-based regression tasks using MLLMs. The application of MLLMs to image-based regression tasks has garnered increasing attention, driven by the potential to leverage the rich representations learned from large-scale multimodal datasets. Image Quality Assessment (IQA), Image Aesthetic Assessment (IAA) and AI-Generated Image Quality Assessment (AIGIQA) are prominent areas where MLLMs have shown promising results. In IQA, the goal is to predict the perceived quality of an image, often by learning to map image features to subjective quality scores. In IAA, the objective is to assess the aesthetic appeal of an image, typically by predicting scores that reflect human aesthetic preferences. In AIGIQA, generated images are evaluated for perceptual quality and alignment with the generation prompt.
One approach to improve IAA representations is through specialized pre-training, as demonstrated by VILA [4]. VILA employs a vision-language pre-training strategy to learn representations that are specifically tailored for aesthetic assessment. This involves training the model on a large dataset of images and associated aesthetic scores, using a combination of contrastive learning and regression objectives. By pre-training on aesthetic-specific data, VILA is able to learn representations that are more effective for predicting aesthetic scores compared to general-purpose vision-language models. The recent Q-Align [20] teaches an MLLM for visual rating aligned with human opinions. Q-Align achieves state-of-the-art performance on image quality assessment (IQA), image aesthetic assessment (IAA), as well as video quality assessment (VQA) tasks. Q-Align unifies the three tasks into one model they call OneAlign.
Architecturally, our work generalizes and extends the Q-Align framework by substituting the vocabulary-based classification head of mPLUG-Owl2 with regression using a classification head, resulting in improved performance and broader applicability. In addition, we present and analyze performance gains that are obtained by the incorporation of textual prompts with semantic relevance to the input images during fine-tuning MLLMs for image-based regression tasks.
A concurrent approach, DeQA-Score [23], addresses the limitations of discretizing continuous quality scores into one-hot labels, which can lead to information loss. DeQA-Score proposes to discretize the entire score distribution into soft labels rather than hard classification targets, aiming to preserve more information about the continuous nature of quality scores. While this approach tackles the discretization problem through distributional modeling, our work demonstrates that simply increasing the number of bins in a straightforward regression using classification framework achieves state-of-the-art performance gains without requiring complex distributional representations (see Fig. 3). This suggests that the discretization accuracy problem can be effectively addressed through increased granularity rather than sophisticated label representations.
@@ 3.1 Architecture overview
RvTC builds upon the Multimodal Large Language Model mPLUG-Owl2 [22], which demonstrates strong visual perception and language understanding capabilities. The base architecture comprises: (1) Vision encoder: ViT-L/14 [15] processes input images. (2) Visual abstractor: Reduces visual features to 64 semantic token embeddings per image. (3) Language decoder: LLaMA-2-7B [18] serves as a universal interface for mixed vision-language input. We replace mPLUG-Owl2’s vocabulary-constrained classification head with a $K$-bin linear classification head that supports arbitrary bin counts for regression tasks. The bin classification head is applied to the penultimate hidden-state embedding of the final token that acts as an aggregator for all of the tokens.
@@ 3.3 Training and inference
For training, we use standard cross-entropy loss to optimize bin classification. During inference, posterior probabilities ${p_{1},p_{2},\ldots,p_{K}}$ are computed via softmax, then converted to continuous values through a weighted sum: $\sum_{i=1}^{K}p_{i}b_{i}$, where $b_{i}$ is the center of bin $i$.
@@ 3.4 Advantages of the bin-based approach
Note that when quantizing the target values of the data, there is freedom in setting the quantization scheme, for example, using a non-uniform quantization scheme. However, increasing the number of bins is a straightforward way to reduce quantization noise and simplify training, removing the need to tune hyperparameters of the range at the cost of added complexity to the classification task.
@@ 3.5 Training configurations
Multimodal training with data-specific prompts. For multimodal training, we incorporate data-specific prompts that contain semantic information relevant to individual images, rather than generic task descriptions. This approach enables the model to leverage cross-modal understanding for improved regression performance.
@@ 4.1 Experimental setup
Datasets. We conduct experiments on five datasets spanning different image assessment tasks. For Image Aesthetic Assessment (IAA), we use AVA [13], a prominent benchmark extracted from the DPChallenge website containing over 250,000 photographic images with mean opinion scores (MOS) from 1-10. Each image belongs to one of 1,400 challenges with descriptive titles such as ”Rule of Thirds” and ”Self Portrait Without People”; these challenge titles serve as our data-specific prompts.
For Image Quality Assessment (IQA), we evaluate on three datasets: KonIQ-10k [2] (10k images from the extensive YFCC100M multimedia database), SPAQ [1] (11k smartphone photos), and KADID-10k [10] (10k images with 25 distortion types at 5 intensity levels). For AI-Generated Image Quality Assessment (AIGIQA), we use AGIQA-3k [9], containing 3k AI-generated images with corresponding generation prompts and MOS ratings for semantic alignment and perceptual quality.
@@
We fine-tune the entire model using Adam optimizer [6] with cosine learning rate scheduling [12]. The learning rate is initialized to 1e-5 with 0.03% warm-up steps. For experiments without textual prompts, we use batch size 128 for 2 epochs; with textual prompts, we train for 3 epochs. All experiments use 4 NVIDIA RTX H100 GPUs.
@@ 4.2 Baseline performance: image-only RvTC
NIMA [16] 0.86/0.90 0.91/0.91 NA/NA
MUSIQ [3] 0.93/0.91 0.92/0.92 NA/NA
LIQE [24] 0.92/0.91 0.92/0.92 0.93/0.93
One-Align [20] 0.94/0.95 0.93/0.93 0.94/0.94
[…]
Comparison with existing methods. Comparing with established baselines (Tab. 1, rows 3-8; Tab. 2), image-only RvTC achieves state-of-the-art performance on AVA with SRCC of 0.833 and PLCC of 0.831, surpassing the previous best method One-Align by 1.0 and 1.2 correlation points respectively. On IQA datasets, RvTC matches or exceeds existing state-of-the-art across all benchmarks: achieving equivalent performance to One-Align on KonIQ-10k and SPAQ, while substantially outperforming all methods on KADID-10k.
@@
[…]
Reversed Syllabus 0.8211 0.8180
Alternative Syllabus 11footnotemark: 1 0.8214 0.8193
Image-Only 0.8229 0.8197
[…]
RvTC - Image-Only (51 bins) 0.8329 0.8314
Impact of regression formulation. The progression from vocabulary-constrained classification to unconstrained bin-classification shows clear benefits. Moving from Q-Align’s 5-token vocabulary approach to RvTC with 5 bins yields similar performance (SRCC 0.823), but increasing to 51 bins provides a substantial boost to 0.833. This demonstrates that discretization granularity through increased bin count is more effective than attempting to align outputs with human vocabulary patterns, eliminating the need for manual vocabulary crafting while achieving superior performance.
@@ 4.3 Impact of data-specific prompts
Performance improvements. Tab. 1 demonstrates the substantial impact of incorporating challenge titles. In the linear probing setting (RvTC-LP+), adding challenge titles improves performance from 0.709 to 0.742 (average SRCC and PLCC), representing a significant gain of 3.3 correlation points without any backbone fine-tuning. This improvement demonstrates that language-based features enhance performance even when the multimodal backbone remains frozen.
The gains become even more pronounced with full fine-tuning. RvTC+ achieves a remarkable correlation of 0.90, improving from our already state-of-the-art image-only baseline of 0.83. This represents a jump of 7 correlation points and establishes a new state-of-the-art on the AVA dataset.
@@ 4.4 Ablation studies and analysis
This section provides ablation studies and analysis to understand the mechanisms behind RvTC’s performance improvements. We establish that a substantial portion of the gains achieved by incorporating challenge titles into RvTC fine-tuning on the AVA dataset stem from cross-modal understanding rather than statistical artifacts. Through controlled experiments, we decompose performance improvements into inter-challenge (driven by statistical biases) and intra-challenge (indicating semantic understanding) components. Additionally, we analyze the impact of bin count on performance across different training configurations, demonstrating that our framework achieves optimal performance through straightforward bin increase. We examine how data-specific prompts interact with our bin-based regression approach, showing that semantic prompts provide stabilizing effects that mitigate overfitting while maintaining performance gains.
@@
Bin count analysis. Fig. 3 (top) demonstrates that performance scales monotonically with bin count across different training lengths, rendering the quantization scheme practically irrelevant once a sufficient number of bins is used. We systematically evaluated 5–101 bins, finding that performance improvement saturates at approximately 51 bins across all datasets. Notably, increasing the number of bins consistently improves correlation scores for all training durations, while longer training has a negative effect, especially for smaller bin counts, due to overconfidence in binning classification. This analysis reveals that optimizing the number of bins in RvTC is straightforward; unlike vocabulary-constrained approaches, our method achieves improved performance and training stability simply by increasing the randomly initialized regression head’s granularity.
Fig. 3 (bottom) reveals how data-specific prompts interact with our framework. Consistent with the image-only setting, increasing the number of bins consistently improves correlation scores across all training durations. However, challenge titles introduce a stabilizing effect: they completely eliminate the negative impact of prolonged training observed in the image-only case when the number of bins is sufficiently large. This suggests that incorporating challenge titles mitigates overfitting in scenarios with fewer bins and prolonged training.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) show promise for image-based regression tasks, but current approaches face key limitations. Recent methods fine-tune MLLMs using preset output vocabularies and generic task-level prompts (e.g., ”How would you rate this image?”), assuming this mimics human rating behavior. Our analysis reveals that these approaches provide no benefit over image-only training. Models using preset vocabularies and generic prompts perform equivalently to image-only models, failing to leverage semantic understanding from textual input. We propose Regression via Transformer-Based Classification (RvTC), which replaces vocabulary-constrained classification with a flexible bin-based approach. Unlike approaches that address discretization errors through complex distributional modeling, RvTC eliminates manual vocabulary crafting through straightforward bin increase, achieving state-of-the-art performance on four image assessment datasets using only images. More importantly, we demonstrate that data-specific prompts dramatically improve performance. Unlike generic task descriptions, prompts containing semantic information about specific images enable MLLMs to leverage cross-modal understanding. On the AVA dataset, adding challenge titles to prompts substantially improves our already state-of-the-art image-only baseline. We demonstrate through empirical evidence from the AVA and AGIQA-3k datasets that MLLMs benefit from semantic prompt information, surpassing mere statistical biases. We validate RvTC across two different MLLM architectures, demonstrating consistent improvements and method generalizability.  [^1] The source code is available at https://github.com/royhj/rvtc.
@@ 1 Introduction
Vision-language models trained on massive unlabeled image-language datasets have demonstrated a remarkable capacity to extract universal image features, with CLIP [16] achieving impressive zero-shot classification performance on benchmark datasets such as ImageNet. Building on these foundations, Multimodal Large Language Models (MLLMs) have evolved to seamlessly fuse image and text embeddings with generative language capabilities [5, 11, 24]. This has sparked growing interest in transferring MLLM capabilities to image-based regression tasks, including Image Quality Assessment (IQA), Image Aesthetics Assessment (IAA) [4, 22], and AI-Generated Image Quality Assessment (AIGIQA) [23, 15].
@@
Table 1: Comparison of models on the AVA dataset. RvTC achieves state-of-the-art correlation of 0.83 without textual input, and rises to 0.90 with data-specific prompts (RvTC+). “LP” denotes linear probing of the regression head; “+” indicates inclusion of challenge titles. Results are reported using mPLUG-Owl2 as backbone (out baseline architecture); see Tab. 6 for validation with QWEN2-VL-2B, which yields even stronger results.
@@
2. Enhancing MLLM performance through data-specific prompts: We demonstrate that generic task prompts (e.g., ”How would you rate this image?”) fail to leverage cross-modal capabilities. In contrast, fine-tuning with data-specific prompts yields substantial improvements in correlation results. For example, the prominent IAA dataset AVA includes challenge titles that characterize images with descriptive phrases such as ”Rule of Thirds” or ”Outdoor Macro Shot.” Incorporating these semantic descriptors during fine-tuning substantially improves correlation results from our already state-of-the-art image-only baseline. We validate this approach across two different MLLM architectures, demonstrating consistent improvements and method generalizability.
@@ 2 Related work
REgression using CLAssification (RECLA). RECLA transforms regression into classification by discretizing continuous targets into bins and training a classifier to predict the appropriate bin [18]. Predictions are mapped back to continuous values through weighted averaging of bin centers. We show that RECLA integrates seamlessly with pre-trained MLLMs through straightforward fine-tuning.
@@
Image assessment regression tasks using MLLMs. Recent work has applied MLLMs such as mPLUG-Owl2 [24] to Image Quality Assessment (IQA), Image Aesthetic Assessment (IAA), and AI-Generated Image Quality Assessment (AIGIQA). In IQA, the goal is to predict the perceived quality of an image, often by learning to map image features to subjective quality scores. In IAA, the objective is to assess the aesthetic appeal of an image, typically by predicting scores that reflect human aesthetic preferences. In AIGIQA, generated images are evaluated for perceptual quality and alignment with the generation prompt. The recent Q-Align [22] teaches an MLLM for visual rating aligned with human opinions. Q-Align achieves state-of-the-art performance on IQA, IAA, and Video Quality Assessment (VQA) tasks. Q-Align unifies the three tasks into one model they call OneAlign.
Our work extends Q-Align by replacing vocabulary-based classification with flexible bin-based regression and demonstrates that semantically relevant textual prompts significantly improve performance. While concurrent work DeQA-Score [25] addresses discretization through soft label distributions, we show that simply increasing bin counts in a straightforward classification framework achieves superior results without complex distributional modeling.
@@ 3.1 Architecture overview
We build our baseline model of RvTC upon the Multimodal Large Language Model mPLUG-Owl2 [24], which demonstrates strong visual perception and language understanding capabilities. The base architecture comprises: (1) Vision encoder: ViT-L/14 [16] processes input images. (2) Visual abstractor: Reduces visual features to 64 semantic token embeddings per image. (3) Language decoder: LLaMA-2-7B [19] serves as a universal interface for mixed vision-language input. We replace mPLUG-Owl2’s vocabulary-constrained classification head with a $K$-bin linear classification head that supports arbitrary bin counts for regression tasks. The bin classification head is applied to the penultimate hidden-state embedding of the final token, which acts as an aggregator for all of the tokens.
@@ 3.3 Training and inference
For training, we use standard cross-entropy loss to optimize bin classification. During inference, posterior probabilities $p_{1},p_{2},\ldots,p_{K}$} are computed via softmax and then converted to continuous values through a weighted sum: $\sum_{i=1}^{K}p_{i}b_{i}$, where $b_{i}$ is the center of bin $i$.
We also evaluated ordinal regression [14] trained on the same K-bin discretization as thresholds for binary cross-entropy. Across all datasets, ordinal regression matched standard cross-entropy within experimental noise and did not yield consistent improvements. For simplicity and reproducibility, we therefore use cross-entropy throughout.
@@ 3.4 Advantages of the bin-based approach
Figure 2: Effect of bin count on quantization noise, measured as (SRCC + PLCC)/2 between AVA ground truth mean opinion score values and their corresponding bin centers
Note that when quantizing the target values of the data, there is freedom in setting the quantization scheme; for example, one may use a non-uniform quantization scheme. However, increasing the number of bins is a straightforward way to reduce quantization noise and simplify training, removing the need to tune hyperparameters of the range at the cost of added complexity to the classification task. Fig. 2 quantifies this effect, demonstrating a near-perfect correlation between original targets and bin centers at 51+ bins.
@@ 3.5 Training configurations
Multimodal training with data-specific prompts. For multimodal training, we incorporate data-specific prompts that contain semantic information relevant to individual images rather than generic task descriptions. This approach, denoted RvTC+, enables the model to leverage cross-modal understanding for improved regression performance.
Throughout this paper, we use RvTC when referring to image-only training and full model fine-tuning; RvTC+ indicates inclusion of data-specific prompts with full fine-tuning, and the suffix ”LP” denotes linear probing, where only the regression head is trained while the backbone remains frozen.
@@ 4.1 Experimental setup
Datasets. We conduct experiments on five datasets spanning different image assessment tasks. For Image Aesthetic Assessment (IAA), we use AVA [13], a prominent benchmark extracted from the DPChallenge website containing over 250,000 photographic images with mean opinion scores (MOS) ranging from 1 to 10. Each image belongs to one of 1,400 challenges with descriptive titles such as ”Rule of Thirds” and ”Self Portrait Without People”; these challenge titles serve as our data-specific prompts.
For Image Quality Assessment (IQA), we evaluate on three datasets: KonIQ-10k [2] (10k images from the extensive YFCC100M multimedia database), SPAQ [1] (11k smartphone photos), and KADID-10k [10] (10k images with 25 distortion types at 5 intensity levels). For AI-Generated Image Quality Assessment (AIGIQA), we use the AGIQA-3k dataset [9], containing 3k AI-generated images with corresponding generation prompts and MOS ratings for semantic alignment and perceptual quality.
@@
We fine-tune the entire model using Adam optimizer [6] with cosine learning rate scheduling [12]. The learning rate is linearly warmed up from 0 to 1e-5 over the first 3% of total training steps. For experiments without textual prompts, we use batch size 128 for 2 epochs; with textual prompts, we train for 3 epochs. All experiments use 4 NVIDIA RTX H100 GPUs.
@@ 4.2 Baseline performance: image-only RvTC
NIMA [17] 0.859/0.896 0.907/0.910 NA/NA
MUSIQ [3] 0.916/0.928 0.917/0.921 NA/NA
LIQE [26] 0.919/0.908 0.922/0.919 0.930/0.931
Q-Align [22] 0.940/0.941 0.930/0.933 0.919/0.918
RvTC (ours) 0.942/0.952 0.926/0.930 0.978/0.981
Table 3: IQA cross-domain results: Ours vs Q-align. Values are SRCC/PLCC, rounded to two decimals. In each cell, the better pair is bolded by the average $(\text{SRCC}+\text{PLCC})/2$.
Train Eval KonIQ SPAQ KADID
KonIQ (ours) 0.942/0.952 0.899/0.897 0.658/0.659
   Q-Align [22] 0.940/0.941 0.887/0.886 0.684/0.674
SPAQ (ours) 0.843/0.874 0.926/0.930 0.677/0.685
   Q-Align [22] 0.848/0.879 0.930/0.933 0.743/0.740
KADID (ours) 0.639/0.657 0.866/0.865 0.978/0.981
   Q-Align [22] 0.668/0.665 0.860/0.854 0.919/0.918
Comparison with existing methods. Comparing with established baselines (Tab. 1, rows 3-8; Tab. 2), image-only RvTC achieves state-of-the-art performance on AVA with SRCC of 0.833 and PLCC of 0.831, surpassing the previous best method One-Align by 1.0 and 1.2 correlation points respectively. On IQA datasets, RvTC matches or exceeds existing state-of-the-art across all benchmarks: achieving equivalent performance to Q-Align on KonIQ-10k and SPAQ, while substantially outperforming all methods on KADID-10k. Furthermore, in the cross-domain evaluation (Tab. 3), RvTC and Q-align achieve broadly similar performance, with each method holding slight advantages in different train–test settings, indicating comparable generalization ability across datasets.
@@
Q-Align (reproduced) 0.821 0.818
Reversed Syllabus 0.821 0.818
Alternative Syllabus 0.821 0.819
Image-Only 0.823 0.820
RvTC - Image-Only (5 bins) 0.823 0.818
RvTC - Image-Only (51 bins) 0.833 0.831
Impact of regression formulation. The progression from vocabulary-constrained classification to unconstrained bin-classification shows clear benefits. Moving from Q-Align’s 5-token vocabulary approach to RvTC with 5 bins yields similar performance (SRCC 0.823), but increasing to 51 bins provides a substantial boost to 0.833. Note that both ”Reversed Syllabus” (which inverts Q-Align’s original vocabulary order) and ”Alternative Syllabus” (using a completely different 5-token vocabulary: ”fish”, ”round”, ”dash”, ”car”, ”end”) achieve nearly identical performance, verifying that results are independent of specific vocabulary choices. This demonstrates that discretization granularity through increased bin count is more effective than attempting to align outputs with human vocabulary patterns, eliminating the need for manual vocabulary crafting while achieving superior performance.
@@ 4.3 Impact of data-specific prompts
Robustness to prompt formulation. To establish the role of prompt design methodology in our approach, we conducted experiments examining the sensitivity of RvTC to prompt template variations. Our analysis on AVA reveals that adding challenge titles without any additional text produces identical results to elaborate prompt templates that incorporate the challenge title, demonstrating that semantic content rather than template structure drives performance. This finding indicates that RvTC’s effectiveness stems from semantic relevance rather than precise prompt engineering, reducing the need for extensive template optimization that typically characterizes other multimodal methods.
[…]
@@ 4.4 Ablation studies and analysis
We conduct ablation studies to understand the mechanisms behind RvTC’s performance improvements. Through controlled experiments, we demonstrate that incorporating challenge titles yields gains from cross-modal understanding rather than statistical artifacts by decomposing improvements into inter-challenge (statistical bias) and intra-challenge (semantic understanding) components. We validate method generalizability across different MLLM architectures and analyze how bin count and data-specific prompts affect performance and training stability.
@@
Table 6: Results across different backbones on the AVA dataset
Base Model RvTC RvTC+
SRCC PLCC SRCC PLCC
mPLUG-Owl2 0.833 0.831 0.899 0.901
QWEN2-VL-2B 0.843 0.842 0.906 0.908
Backbone generalization. To address the generalizability of our RvTC method across MLLMs, we reproduce our experiments on AVA using QWEN2-VL-2B [20], a recent vision-language model with different architectural characteristics than mPLUG-Owl2. We use identical training procedures and hyperparameters as described in Section 4.1. The improvement from RvTC to RvTC+ on both models suggests that our method captures fundamental aspects of vision-language reasoning rather than exploiting model-specific architectural biases. This validates that data-specific prompts enable effective cross-modal understanding in MLLMs in fine-tuning for image-based regression tasks.
@@
Bin count analysis. Fig. 4 demonstrates that performance scales monotonically with bin count across different training lengths, with improvement saturating at approximately 51 bins. While increasing bins consistently improves correlation scores, longer training degrades performance in image-only settings due to overconfidence in binning classification.
However, data-specific prompts change this dynamic. Fig. 4 (bottom) shows that challenge titles completely eliminate the negative impact of prolonged training when sufficient bins are used, introducing a stabilizing effect that mitigates overfitting scenarios.
Training stability advantage. This stabilization represents a key practical advantage over vocabulary-constrained approaches. While previous methods require careful training duration tuning to avoid degradation, RvTC achieves robust stability through increased bin granularity. This combination enables longer optimization without extensive hyperparameter tuning, making RvTC more practical for deployment than existing multimodal regression methods.